A steel rod of diameter 14cm subjected to a load of 50N,with a strain of 20. Calculate the young modulus or modulus of elasticity.
1
Expert's answer
2020-09-02T12:14:32-0400

The stress

"\sigma=\frac{F}{A}=\varepsilon E"

Hence, the modulus of elasticity

"E=\frac{F}{\varepsilon A}=\rm\frac{50\: N}{20\times 3.14\times (0.07\: m)^2}=162\:Pa"